ALTOGETHER 568 TICKETS WERE SOLD TO A PLAY. AN ADULT TICKET COSTS $5.00. A CHILD TICKET COSTS $2.OO. TICKET SALES WERE $2345.00. THE THEATER MANAGER EXPECTS TO SELL THE SAME NUMBER OF TICKETS TO THE NEXT PLAY, BUT WANTS TO INCREASE THE TICKET SALES TO $3554 BY INCREASING THE PRICE OF ADULT TICKETS. FIND THE NEW PRICE OF AN ADULT TICKET.
Answer by ankor@dixie-net.com(22740)

Two equations:
a + c = 568
Rearrange to use for substitution, subtract c from both sides:
a = (568 - c)
:
5a + 2c = 2345
:
Substitute (568-c) for a;
5(568-c) + 2c = 2345
2840 - 5c + 2c = 2345
-3c = 2345 - 2840
-3c = -495
c = -495/-3
c = +165 children
:
Find the adults using equation: a + c = 568
a + 165 = 568
a = 568 - 165
a = 403 adults
:
Check solutions in the equation 5a + 2c = 2345
5(403) + 2(165) =
2015 + 330 = 2345
:
Let p = the new adult price:
403p + 165(2) = 3554
403p + 330 = 3554
403p = 3554 - 330
403p = 3224
p = 3224/403
p = $8 is the new price for adults